NAOJ GW Elog Logbook 3.2
Displaying report 1-1 of 1.
AkihiroTomura - 11:07, Wednesday 31 January 2018 (647)

Comment to SHG characterization (Click here to view original report: 627)
I did the SHG output power vs input power measurement again. I used HWP and FI pair after the main laser to change the input power in order not to get mode-hops. During the measurement, the current value of the main laser was 2.004 A. The SHG crystal temerature was 3.170 kOhm.
Images attached to this comment